OpenStudy (anonymous):

OK can someone please check my answer for me the problem is x^2-2x-1=0 it states that it has 2 solutions it wants me to list the smallest solution which i calculated to me 0 but the software marks that wrong is there something i did wrong?

jimthompson5910 (jim_thompson5910):

What 2 solutions did you get?

OpenStudy (anonymous):

um i believe I got 0 and i forgot what the other was i don't have the work all together right now but it was bigger then 0

OpenStudy (anonymous):

i know that for sure

jimthompson5910 (jim_thompson5910):

0 isn't a solution x^2-2x-1=0 (0)^2-2(0)-1=0 ... plug in x = 0 0 - 2(0) - 1 = 0 0 - 0 - 1 = 0 -1 = 0 ... this is false, so x = 0 isn't a solution

OpenStudy (anonymous):

ok i see my error but can you help me solve the problem then because i don't know what I'm really doing

jimthompson5910 (jim_thompson5910):

You have to use the quadratic formula to solve for x.

jimthompson5910 (jim_thompson5910):

Do you remember that formula?

OpenStudy (anonymous):

yes i remember it

OpenStudy (anonymous):

thats what I'm having the issue with someone helped me get the numbers i need to plug in but its still confusing me

OpenStudy (javk):

It would be difficult to solve using factorization so use the formula \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

ok thats what I used originally I was helped to get the numbers to plug in for the variables we got -2 for b 1 for a -1 for c

OpenStudy (anonymous):

but somewhere during the solving process it went wrong i don't really know what i did wrong

jimthompson5910 (jim_thompson5910):

what is \(\Large b ^{2}-4ac\) equal to?

OpenStudy (anonymous):

that would be -2^2 which would be -4 and -4(1)(-1) would be-4+4 which would be 0 right?

jimthompson5910 (jim_thompson5910):

b^2 is not -2^2 = -4 b^2 = (-2)^2 = 4

OpenStudy (anonymous):

ok sorry your right i don't know what i was thinking its getting late i just want to hurry and finish this last problem

jimthompson5910 (jim_thompson5910):

so what is b^2 - 4ac equal to?

OpenStudy (anonymous):

ok so b^2 = 4 and -4(1)(-1)= 4 so its 4+4 which equals 8 right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

\[\large x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] \[\large x=\frac{ -(-2) \pm \sqrt{(-2) ^{2}-4(1)(-1)} }{ 2(1) }\] \[\large x=\frac{ 2 \pm \sqrt{8} }{ 2 }\] how do we simplify sqrt(8) ?

OpenStudy (anonymous):

it can be broken down to 2*2*2 and you can take 2 2s so it would then be 2 square root 2 right?

jimthompson5910 (jim_thompson5910):

yes \[\large x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] \[\large x=\frac{ -(-2) \pm \sqrt{(-2) ^{2}-4(1)(-1)} }{ 2(1) }\] \[\large x=\frac{ 2 \pm \sqrt{8} }{ 2 }\] \[\large x=\frac{ 2 \pm 2*\sqrt{2} }{ 2 }\]

jimthompson5910 (jim_thompson5910):

then we can factor out 2 from the numerator and cancel the '2's \[\large x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] \[\large x=\frac{ -(-2) \pm \sqrt{(-2) ^{2}-4(1)(-1)} }{ 2(1) }\] \[\large x=\frac{ 2 \pm \sqrt{8} }{ 2 }\] \[\large x=\frac{ 2 \pm 2*\sqrt{2} }{ 2 }\] \[\large x=\frac{ 2(1 \pm \sqrt{2}) }{ 2 }\] \[\large x=\frac{ \cancel{2}(1 \pm \sqrt{2}) }{ \cancel{2} }\] \[\large x=1 \pm \sqrt{2}\] \[\large x=1 + \sqrt{2} \ \ \text{or} \ \ x = 1-\sqrt{2}\]

OpenStudy (anonymous):

ok so x= 1-square root of 2 would be my answer since its asking for the smallest solution?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok thanks for your help

OpenStudy (anonymous):

now to go sleep lol so i can start thinking straight again